# “I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

### v1

@@ 3.3.1. Preparation Phase
Prior work has shown that people tend to view an AI system as more fair when they personally receive a favorable outcome (Wang et al., 2020), for example, a “Good Credit” decision in our credit rating scenario. To ensure an objective assessment and remain unaffected by favorable or unfavorable AI outcomes, our participants, as decision subjects, were not given any specific decision label (i.e., “Good Credit” or “Bad Credit”).
@@ 3.4.1. Functionalities Supporting Study Tasks
We designed the prototype system to support participants in the Main Phase tasks. Building on prior work that highlights the value of contextual information for fairness-related decision making (Cheng et al., 2021; Ashktorab et al., 2023), our prototype provided such information with metric explanations to help participants reflect on fairness. We outline its core functionalities, with technical details in Appendix A.3.
Task 1: Feature Selection & Ranking To support participants in completing Task 1, the system offered the following functionalities:
  • Participants could use the prototype system to explore the cases, the model, and the features used for the model training (Figure 1). They were able to investigate each case’s feature values along with its ground-truth label and the AI’s predicted outcome (Figure 1 A).
@@
  • Inspired by prior work showing that non-expert stakeholders benefit from transparency to support comprehension and fairness reasoning (Yan et al., 2020; Nakao et al., 2023, 2022), participants could explore feature information (Figure 1 C) through causal graphs showing feature relationships in the training and test data (generated with CausalNex (Zheng et al., 2018)); model-based feature importance, and training data distributions. These elements acted as “white-box” explanations, exposing the model’s inner workings and training context (Cheng et al., 2019). Importantly, in our settings, these visualizations were not intended to steer participants toward particular features but to situate them in context. We also reminded participants that such information may itself reflect biases in the dataset, model, or labels, underscoring its role as transparency rather than prescriptive guidance. Participants were free to use or disregard them, combining system-provided insights with their own experience.
@@
Figure 1. Prototype System: Dashboard (also, Task 1). Component A presents the test dataset in a tabular format, displaying each case’s feature values, its ground-truth label, and the AI’s predicted outcome. Component B provides AI model explanations, performance, and information about protected features. Component C contains four sub-components: Causal Graph: Training Data and Causal Graph: Test Data, which visualize feature relationships in the training and test sets, allowing participants to click on a feature to highlight its connected edges; Feature Importance highlights the AI model’s key influencing features, and Data Distribution displays the feature distribution within the training set. Component D allows participants to select and rank features by using checkboxes and dragging them into their preferred order. Clicking Confirm Feature Selection & Ranking navigates to Task 2.
@@
Figure 2. Prototype System: Feature-Metric Pairing Module (Task 2). Component E shows a participant viewing (Group Fairness - Equal Opportunity) for Gender. The interface provides a lay definition, a bar chart with metric results, and metric calculation. Next, it shows a transparent individual-level visualization of the metric calculation, where applicants are sub-grouped by gender and AI prediction (i.e., “Predicted Credit is Good” and “Predicted Credit is Bad”), and then further grouped by ground-truth label (“Real Credit is Good” and “Real Credit is Bad”). The gray region denotes applicants with real bad credit, and are excluded from this metric calculation. For applicants with real good credit, the metric is calculated as the proportion of Yellow (real good & predicted good) over Yellow + White (all real good).
Unlike prior systems (Luo et al., 2025; Cheng et al., 2021) which were designed around AI experts’ norms for assessing fairness, our prototype supports 1) participants to apply fairness metrics to any feature they considered relevant and 2) to help make metrics accessible and understandable. Beyond technical definitions, we also included descriptions of each metric’s strengths and limitations, informed by feedback from non-experts (Luo et al., 2025; Cheng et al., 2021). These descriptions aimed to reduce the cognitive burden of metric choice, make selection more contextual, and reflect non-experts’ preference for understanding the practical implications of fairness metrics (Luo et al., 2025), thereby supporting participants in aligning metrics with their own values. Further, we introduced a “Custom Fairness” option that allowed participants to define their own metrics when predefined metrics did not capture their fairness concerns. Given the lack of prior work on non-experts’ custom metric design, we implemented this as an open-ended form: participants verbally expressed definitions while researchers documented and probed for clarification; participants could also handwrite their definitions if preferred.
@@ 3.4.2. Comparison with Existing Tools
We detailed the design requirements identified from people without AI expertise (Nakao et al., 2023), and compared our prototype system with other common tools in fairness assessments also designed for non-expert people: FET (Cheng et al., 2021), FairHIL (Nakao et al., 2023), FEE (Luo et al., 2025), highlighting how each tool addressed the needs of participants without AI expertise (Table 1). Compared to existing tools, our prototype addresses multiple design requirements and further extends functionality through feature selection, feature-metric pairing, and priority-based ranking.
@@ 4. Results
4.1. How do stakeholders identify features that are crucial for assessing AI fairness, and how do they prioritize them? (RQ1)
Our study uncovered that selecting features for fairness assessment of AI systems involved complex decision-making for participants. We also explored the underlying reasons for their choices through a thematic analysis and the themes that were extracted are indicated in bold text.
4.1.1. Divergent Fairness Perspectives in Feature Selection
@@
We observed that participants had different perspectives on AI fairness, which influenced how they chose the features they considered important for evaluating whether an AI system is fair. Only a few participants (4 participants) held the perspective of affordability fairness (Nakao et al., 2022), aiming for fairness by basing AI decisions on an individual’s actual ability to afford a loan. As P13 emphasized, “I will care more about this applicant’s creditness. Um. This includes […] like some features to see whether they can pay back or should get the loan or not”. Similarly, P5 mentioned, “Because the main objective of the bank is not just to do the things on the basis of morality only”. Therefore, this perspective led them to select highly context-specific features that are considered reasonable and precise for evaluating an individual’s credit level, like ‘Credit Amount’, ‘Existing Credits’, and ‘Checking Account’, while deliberately disregarding demographic characteristics such as ‘Gender’, which seem unrelated to credit evaluation but can introduce bias into the prediction. Among these four participants, two participants highlighted the importance of procedural fairness, emphasizing that all individuals should undergo the same evaluation process when assessing affordability. As P5 explained, “If all the persons go from the same litmus paper, [using] same features, same characteristics, so it is, I think, fairer than fair. It is fair and transparent”.
@@
In contrast, the most commonly chosen fairness perspective was outcome fairness (26 participants). Participants with this perspective selected features that they “don’t want this AI to discriminate against” (P30), such as ‘Gender’, which could disproportionately harm or benefit certain individuals or groups. At the same time, they avoided selecting those highly context-specific features, like the previously mentioned ‘Credit Amount’, which are considered reasonable and precise for evaluating an individual’s affordability or credit level. Additionally, three participants among them mentioned a stricter fairness through unawareness fairness notion (Kusner et al., 2017; Hort et al., 2024), explicitly expressing their preference for excluding features such as ‘Telephone’, ‘Gender’, ‘Foreign Worker’, and ‘Residence Length’ when training AI, as they viewed these features as irrelevant to the AI’s decisions.
The remainder of our analysis focuses on the 26 participants who adopted the outcome fairness perspective. Accordingly, all subsequent analyses, including feature selection, metric selection, and threshold setting, are drawn from these 26 participants. The full results of these 26 participants are shown in Appendix A.6. We excluded data from participants who held the affordability fairness perspective, as well as the data related to the features excluded under the fairness through unawareness notion by participants who held an outcome fairness perspective because the metrics we offered were not compatible with these views.
@@
[Uncaptioned image] Figure 3. Top 6 most selected features and their selection frequencies in the Top 3 ranking positions.     [Uncaptioned image] Figure 4. Top 6 most selected features and the distribution of their paired fairness metrics.
@@ 4.1.2. Broad Fairness Considerations
We found that participants demonstrated broader fairness considerations in outcome fairness. In the credit rating scenario, AI experts often select only three features: two protected features (‘Age’ and ‘Gender’) and one sensitive feature (‘Foreign Worker’) (Le Quy et al., 2022). In contrast, 19 of the 20 available features were selected at least once by participants, except for ‘Credit Amount’. In average, our participants selected six features for this credit rating scenario’s fairness assessment (Mean = 6.15, Min = 1, Max = 11, Mode = 6, SD = 2.30), with the most frequently selected features overall being ‘Gender’ (21 participants), ‘Foreign Worker’ (21 participants), ‘Age’ (20 participants), ‘Telephone’ (15 participants), ‘Dependents’ (12 participants), and ‘Purpose’ (10 participants). These broad choices underscore the wide variation in how stakeholders interpret fairness, extending far beyond the narrow sets of protected features typically considered in technical fairness.
@@ 4.1.3. Protected Features’ Role: Central Yet Contested
Our analysis showed that protected and sensitive features played a major role in participants’ fairness assessments, but participants did not agree on their priority for fairness assessment. ‘Gender’, ‘Foreign Worker’, and ‘Age’ were the top three most frequently selected features overall, and these three features also dominated the top three ranking positions (i.e., Top-1, Top-2, and Top-3), as shown in Figure 4, with ‘Gender’ chosen by 18 participants mainly in the Top-1 position, ‘Age’ by 13 participants mainly in the Top-2 position, and ‘Foreign Worker’ by 11 participants mainly in the Top-2 and Top-3 positions. However, not all participants prioritized these protected and sensitive features in their fairness assessment. Only 18 out of 26 participants placed any of these features in their Top-1 position. Additionally, not everyone included any of these features in their selection at all, for example, ‘Gender’ and ‘Foreign Worker’ were omitted by 5 participants, and ‘Age’ was excluded by 4 participants.
@@ 4.1.4. Unprotected Features’ Role: Complementary Yet Indispensable
Our analysis also showed that non-protected features play a complementary yet indispensable role in participants’ fairness assessments, a dimension often overlooked from a technical fairness perspective. Nearly all participants (25 out of 26) selected a combination of both protected and non-protected features to create a more holistic fairness assessment. Even though they are ranked generally lower than protected and sensitive features, nonetheless, they appeared frequently in the Top-3 position. As Figure 4 shows, ‘Telephone’ was ranked highly by 8 participants, sometimes even appearing in Top-1 position, ‘Purpose’ by 8 participants often appearing in the Top-3 position, and ‘Dependents’ by 4 participants. Many non-protected features, such as, ‘Job’, ‘Credit History’, were also selected but ranked much lower, associated with context-specific reasons such as Financial ability considerations (19 participants).
@@ 4.1.5. Shifting Priorities: The Dynamic Nature of Fairness Decisions
Findings summary: Participants demonstrated diverse and complex perspectives in fairness assessment, with the majority basing their feature selections and rankings on outcome fairness. They tended to go beyond legally protected features and considered the potential biases introduced by non-protected features used in AI’s decision-making. This meant more features were attended to, and thus prioritized. Prioritizations were not static; participants changed how they ranked features as they engaged more deeply in fairness assessments.
@@
We investigated the overall metric preferences based on the eight predefined metrics and the ‘Custom Fairness’ metric option. The three most chosen metrics were Counterfactual Fairness, selected 31 times across all participants, Custom Fairness (28 times) and Equal Opportunity (20 times). Less frequently selected metrics were Conditional Statistical Parity (18 times), Consistency (18 times), and Equalized Odds (10 times). This highlights that participants’ fairness needs were diverse, spanning both individual and group fairness categories. Participants also appeared to frequently select more complex metrics such as counterfactual fairness and custom fairness, citing the necessity of assessing individual circumstances and the requirement for flexible defining fairness. However, AI experts often choose metrics that are easy to implement and are often used in legal settings, such as Demographic Parity - further highlighting a misalignment between current technical fairness and human-centered fairness.

### v2

@@ 3.3.1. Preparation Phase
Prior work has shown that people tend to view an AI system as more fair when they personally receive a favorable outcome (Wang et al., 2020), for example, a “Good Credit” decision in our credit rating scenario. To ensure an objective assessment and remain unaffected by favorable or unfavorable AI outcomes, our participants, as decision subjects, were not told that any specific decision label applied to themselves (i.e., whether they would personally receive “Good Credit” or “Bad Credit”) when assessing fairness.
@@ 3.3.3. Post-Task Phase
Figure 1. Prototype System: Dashboard (also, Task 1). Component A presents the test dataset in a tabular format, where each row represents one case (i.e., applicant) of the test dataset, displaying each applicant’s feature values, the ground-truth label, and the AI’s predicted outcome. Component B provides AI model explanations, performance, and information about protected features. Component C contains four sub-components: Causal Graph: Training Data and Causal Graph: Test Data, which visualize feature relationships in the training and test sets, allowing participants to click on a feature to highlight its connected edges; Feature Importance highlights the AI model’s key influencing features, and Data Distribution displays the feature distribution within the training set. Component D allows participants to select and rank features by using checkboxes and dragging them into their preferred order. Clicking Confirm Feature Selection & Ranking navigates to Task 2.
Prototype System Dashboard Prototype system dashboard used in Task 1: feature selection and ranking. It shows the test dataset with feature values, model predictions, and explanations. Visualizations illustrate feature relationships, feature importance, and data distributions. The interface also lets participants select and rank features before proceeding to the next task.
Figure 2. Prototype System: Feature-Metric Pairing Module (Task 2). Component E shows a participant viewing (Group Fairness - Equal Opportunity) for Gender. The interface provides a lay definition, a bar chart with metric results, and metric calculation. Next, it shows a transparent individual-level visualization of the metric calculation, where the test dataset–all applicants–are sub-grouped by gender and AI prediction (i.e., “Predicted Credit is Good” and “Predicted Credit is Bad”), and then further grouped by ground-truth label (“Real Credit is Good” and “Real Credit is Bad”). The gray region denotes applicants with real bad credit, and are excluded from this metric calculation. For applicants with real good credit, the metric is calculated as the proportion of Yellow (real good & predicted good) over Yellow + White (all real good).
Prototype System Metric Selection Prototype system module for Task 2: metric selection. It lets participants pair a fairness metric with a feature, then shows the metric’s explanations, visualizations, and detailed calculations.
@@ 3.4.1. Functionalities Supporting Study Tasks
We designed the prototype system to support participants in the Main Phase tasks to assess AI outcome fairness. Building on prior work that highlights the value of contextual information for fairness-related decision making (Cheng et al., 2021; Ashktorab et al., 2023), our prototype provided such information with metric explanations to help participants reflect on fairness. We outline its core functionalities, with technical details in Appendix A.4, and make the source code publicly available on GitHub.  [^1] The prototype source code is publicly available on GitHub: https://github.com/LInnnLUooo/Lay-Stakeholder-AI-Fairness-Assessment-Tool
Task 1: Feature Selection & Ranking To support participants in completing Task 1, the system offered following functionalities:
  • Participants could use the prototype system to explore each case (i.e., each applicant) from the test dataset, the model, and the features with their descriptions used for the model training (Figure 1). They were able to investigate each applicant’s feature values along with their ground-truth label and the AI’s predicted outcome (Figure 1 A).
@@
  • Inspired by prior work showing that lay stakeholders benefit from transparency to support comprehension and fairness reasoning (Yan et al., 2020; Nakao et al., 2023, 2022), participants could explore feature information (Figure 1 C) through causal graphs showing feature relationships in the training and test data (generated with CausalNex (Zheng et al., 2018)); model-based feature importance, and training data distributions. These elements acted as “white-box” explanations, exposing the model’s inner workings and training context (Cheng et al., 2019). Importantly, in our settings, these visualizations were not intended to steer participants toward particular features but to situate them in context. We also reminded participants that such information may itself reflect biases in the dataset, model, or labels, underscoring its role as transparency rather than prescriptive guidance. Participants were free to use or disregard them, combining system-provided insights with their own experience.
@@
Unlike prior systems (Luo et al., 2025; Cheng et al., 2021) which were designed around AI experts’ norms for assessing fairness, our prototype supports 1) participants to apply fairness metrics to any feature they considered relevant and 2) to help make metrics accessible and understandable. Beyond technical definitions, we also included descriptions of each metric’s strengths and limitations, informed by feedback from lay stakeholders (Luo et al., 2025; Cheng et al., 2021). These descriptions aimed to reduce the cognitive burden of metric choice, make selection more contextual, and reflect lay stakeholders’ preference for understanding the practical implications of fairness metrics (Luo et al., 2025), thereby supporting participants in aligning metrics with their own values. Further, we introduced a “Custom Fairness” option that allowed participants to define their own metrics when predefined metrics did not capture their fairness concerns. Given the lack of prior work on lay stakeholders’ custom metric design, we implemented this as an open-ended form: participants verbally expressed definitions while researchers documented and probed for clarification; participants could also handwrite their definitions if preferred.
@@ 3.4.2. Comparison with Existing Tools
We detailed the design requirements identified from people without AI expertise (Nakao et al., 2023), and compared our prototype system with other common tools in fairness assessments also designed for lay stakeholders: FET (Cheng et al., 2021), FairHIL (Nakao et al., 2023), FEE (Luo et al., 2025), highlighting how each tool addressed the needs of participants without AI expertise (Table 1). Compared to existing tools, our prototype addresses multiple design requirements and further extends functionality through feature selection, feature-metric pairing, and priority-based ranking.
@@ 3.5. Data Collection and Analysis
Because our research specifically focuses on the outcome fairness perspective, our data analysis draws on data from 26 out of 30 participants who assessed fairness through this perspective. To maintain conceptual clarity and a coherent outcome fairness assessment, we excluded all collected data that did not reflect an outcome fairness assessment, i.e., complete data from the remaining 4 participants who held the affordability fairness perspective (Nakao et al., 2022), and the partial data reflecting the fairness through unawareness notion (Kusner et al., 2017; Hort et al., 2024) from 3 of the 26 participants who otherwise held an outcome fairness perspective. These excluded perspectives focused on feature usage during ‘model training’: caring about which features should be included (affordability fairness) or excluded (fairness through unawareness). When participants assessed fairness through these perspectives, they were not evaluating whether the model’s prediction outcomes were biased with respect to those features, nor were they engaging with answering: whom to protect, how to protect, and to what extent. Although we noted that participants with different perspectives sometimes raised the same features (e.g., Gender), indicating concern about the same type of bias, they were reasoning and evaluating fairness from fundamentally different stages of the model pipeline. Table 5 in Appendix A.2 provides an overview of participants whose data were included or excluded.
@@ 4. Results
We first separately report the fairness perspectives participants expressed other than outcome fairness for completeness. We then answer our research questions (RQ1 and RQ2) to uncover how participants assessed AI outcome fairness. Finally, we describe participants’ user experience during the assessment, including the challenges and benefits they encountered and the prototype system’s SUS score.
4.1. Other Fairness Perspectives
@@
Only a few participants (4 participants; P5, P8, P13, P14) held the perspective of affordability fairness (Nakao et al., 2022), aiming for fairness by basing AI decisions on an individual’s actual ability to afford a loan. As P13 emphasized, “I will care more about this applicant’s creditness. Um. This includes […] like some features to see whether they can pay back or should get the loan or not”. Similarly, P5 mentioned, “Because the main objective of the bank is not just to do the things on the basis of morality only”. Therefore, this perspective led them to select highly context-specific features that are considered reasonable and precise for evaluating an individual’s credit level, like ‘Credit Amount’, ‘Existing Credits’, and ‘Checking Account’, while deliberately disregarding demographic characteristics such as ‘Gender’, which seem unrelated to credit evaluation but can introduce bias into the prediction. Among these four participants, two participants highlighted the importance of procedural fairness, emphasizing that all individuals should undergo the same evaluation process when assessing affordability. As P5 explained, “If all the persons go from the same litmus paper, [using] same features, same characteristics, so it is, I think, fairer than fair. It is fair and transparent”. Taken together, these four participants approached fairness by reasoning about how the AI model should be trained, that is, which features it ought to use to accurately capture people’s affordability, rather than, in terms of whether the AI system’s outcomes were fair for the groups or features they sought to protect.
@@
Additionally, three participants (P9, P22, P25) among the 26 participants who held an outcome fairness perspective also mentioned fairness through unawareness fairness notion (Kusner et al., 2017; Hort et al., 2024), explicitly expressing their preference for excluding features such as ‘Telephone’, ‘Gender’, ‘Foreign Worker’, and ‘Residence Length’ when training the AI model, as they viewed these features as irrelevant to the AI’s decisions.
4.2. How do stakeholders identify features that are crucial for assessing AI outcome fairness, and how do they prioritize them? (RQ1)
Our study uncovered that selecting features for fairness assessment of AI systems involved complex decision-making for participants. We also explored the underlying reasons for their choices through a thematic analysis and the themes that were extracted are indicated in bold text. The remainder of our analysis focuses on the 26 participants who assessed outcome fairness. Accordingly, all subsequent analyses, including feature selection and ranking, metric selection, and threshold setting, are drawn from these 26 participants. The full results of these 26 participants are shown in Appendix A.7.
@@
(a) Top 6 most selected features and their selection frequencies in the Top 3 ranking positions.
Heatmap showing the six most frequently selected features and how often they appeared in participants’ top three rankings.
(b) Top 6 most selected features and the distribution of their paired fairness metrics.
Heatmap showing the six most frequently selected features and the distribution of fairness metrics paired with each feature.
Figure 3. Heatmaps: Distribution of participants’ top-selected features and their associated fairness metrics.
The figure on the left is a heatmap showing the six most frequently selected features and how often they appeared in participants’ top three rankings. The figure on the right is a heatmap showing the six most frequently selected features and the distribution of fairness metrics paired with each feature.
@@ 4.2.1. Broad Fairness Considerations
We observed that participants, when assessing outcome fairness, selected features that associated with whom they want to protect, i.e., features they “don’t want this AI to discriminate against” (P30), such as ‘Gender’, which could disproportionately harm or benefit certain individuals or groups. At the same time, they avoided selecting those highly context-specific features, such as ‘Credit Amount’, which they considered reasonable and precise for producing prediction outcomes.
We found that participants demonstrated broader fairness considerations in outcome fairness than AI experts typically do. Previous work (Le Quy et al., 2022) has shown that for the German Credit Dataset, AI experts often select only three features: two protected features (‘Age’ and ‘Gender’) and one sensitive feature (‘Foreign Worker’). In other credit scenarios (de Castro Vieira et al., 2025), AI experts’ attention similarly tends to remain on a narrow set of protected or sensitive features (typically centered on gender and race), with relatively little attention paid to other potential sources of bias. In contrast, 19 of the 20 available features were selected at least once by participants, except for ‘Credit Amount’. In average, our participants selected six features for this credit rating scenario’s fairness assessment (Mean = 6.15, Min = 1, Max = 11, Mode = 6, SD = 2.30), with the most frequently selected features overall being ‘Gender’ (21 participants), ‘Foreign Worker’ (21 participants), ‘Age’ (20 participants), ‘Telephone’ (15 participants), ‘Dependents’ (12 participants), and ‘Purpose’ (10 participants). These broad choices underscore the wide variation in how stakeholders interpret fairness, extending far beyond the narrow sets of protected features typically considered in technical fairness (de Castro Vieira et al., 2025; Le Quy et al., 2022).
@@ 4.2.2. Protected Features’ Role: Central Yet Contested
Our analysis showed that protected and sensitive features played a major role in participants’ fairness assessments, but participants did not agree on their priority for fairness assessment. ‘Gender’, ‘Foreign Worker’, and ‘Age’ were the top three most frequently selected features overall, and these three features also dominated the top three ranking positions (i.e., Top-1, Top-2, and Top-3), as shown in Figure 3(a), with ‘Gender’ chosen by 18 participants mainly in the Top-1 position, ‘Age’ by 13 participants mainly in the Top-2 position, and ‘Foreign Worker’ by 11 participants mainly in the Top-2 and Top-3 positions. However, not all participants prioritized these protected and sensitive features in their fairness assessment. Only 18 out of 26 participants placed any of these features in their Top-1 position. Additionally, not everyone included any of these features in their selection at all, for example, ‘Gender’ and ‘Foreign Worker’ were omitted by 5 participants, and ‘Age’ was excluded by 4 participants.
@@ 4.2.3. Unprotected Features’ Role: Complementary Yet Indispensable
Our analysis also showed that non-protected features play a complementary yet indispensable role in participants’ fairness assessments, a dimension often overlooked from a technical fairness perspective by AI experts (de Castro Vieira et al., 2025; Le Quy et al., 2022). Nearly all participants (25 out of 26) selected a combination of both protected and non-protected features to create a more holistic fairness assessment. Even though they are ranked generally lower than protected and sensitive features, nonetheless, they appeared frequently in the Top-3 position. As Figure 3(a) shows, ‘Telephone’ was ranked highly by 8 participants, sometimes even appearing in Top-1 position, ‘Purpose’ by 8 participants often appearing in the Top-3 position, and ‘Dependents’ by 4 participants. Many non-protected features, such as, ‘Job’, ‘Credit History’, were also selected but ranked much lower, associated with context-specific reasons such as Financial ability considerations (19 participants).
@@ 4.2.4. Shifting Priorities: The Dynamic Nature of Fairness Decisions
Findings summary: Participants demonstrated broad and complex fairness concerns when selecting features to assess whether the model’s outcomes discriminate based on those features. They tended to go beyond legally protected features and considered the potential biases introduced by non-protected features. This meant more features were attended to, and thus prioritized. Prioritizations were not static; participants changed how they ranked features as they engaged more deeply in fairness assessments.
@@
We investigated the overall metric preferences based on the eight predefined metrics and the ‘Custom Fairness’ metric option. The three most chosen metrics were Counterfactual Fairness, selected 31 times across all participants, Custom Fairness (28 times) and Equal Opportunity (20 times). Less frequently selected metrics were Conditional Statistical Parity (18 times), Consistency (18 times), and Equalized Odds (10 times). This highlights that participants’ fairness needs were diverse, spanning both individual and group fairness categories. Participants also appeared to frequently select more nuanced and complex metrics such as counterfactual fairness and custom fairness, citing the necessity of assessing individual circumstances and the requirement for flexible defining fairness.
In contrast, AI experts often choose group fairness metrics that are easier to implement and widely used in legal or compliance settings (Binns, 2020; Selbst et al., 2019), such as Demographic Parity. Prior work commonly recognizes Demographic Parity as one of the most frequently used fairness metrics, both within credit-related contexts (de Castro Vieira et al., 2025) and across other AI applications scenarios (Mehrabi et al., 2021; Verma and Rubin, 2018; Hort et al., 2024). However, this metric was not selected by any of our participants. Similarly, individual fairness metrics such as Counterfactual Fairness and Consistency, as well as more nuanced and more complex-to-implement group fairness metrics like Conditional Statistical Parity, were popular among participants but appear to be used less often by AI experts (de Castro Vieira et al., 2025; Hort et al., 2024; Mehrabi et al., 2021; Verma and Rubin, 2018). This difference in metric preferences highlights a potential misalignment between current technical fairness practices and human-centered fairness expectations.